Problem #7: Use the chain rule to find the value of op when t= 1. z = 4x2 – 5y, x = P +3, y = 2t – 2 Problem #7:
Answer #1

z=4x^2-5y, \ x=t^3+3,\ y=2t-2.\\ \frac{dz}{dt}=\frac{d}{dt}(4x^2-5y)=4.\frac{d}{dt}(x^2)-5.\frac{d}{dt}(y)=4.\frac{d}{dx}(x^2).\frac{dx}{dt}-5.\frac{dy}{dt},\ [using\ chain\ rule]\\ =4.(2x).\frac{d}{dt}(t^3+3)-5.\frac{d}{dt}(2t-2)=8.(t^3+3).3t^2-10.\\ at\ t=1, \frac{dz}{dt}=8.(1+3).3-10=86.(ans)